OpenStudy (shamil98):

How many triples of positive integers (a,b,c) satisfy: \[\ a \le b \le c \] when \[\ abc = 1,000,000,000\]

OpenStudy (anonymous):

Trial and error kind of thing?

OpenStudy (shamil98):

It's combinatorics. I have no idea how to go about this one.

OpenStudy (anonymous):

Wow. Cool problem! I'll have to think about this. Hold on.

OpenStudy (anonymous):

To start, the prime factorization of our number is 2^9 * 5^9 All of our various possibilities will have those 2's and 5's assigned in various ways to a, b and c. Note that it's ok for none of these to be assigned to one two of the variables as we can just make that variable be 1.

OpenStudy (anonymous):

Prime factorization is important here.

OpenStudy (anonymous):

Since \[ 10^9 = 2^95^9 \]Then it is just a matter of dividing up \(2\)s and \(5\)s among \(a\), \(b\), and \(c\). Normally any partition we created could be reused by swapping one variable for another, but since \(a\leq b\leq c\), we must sort our partitions so order does not matter.

OpenStudy (shamil98):

Hm, so how long do you think it would take to figure out all the possible combinations? Are there any other methods?

OpenStudy (anonymous):

There are smart ways of calculating the partitions.